The following narrow-sense heritability values wer
Answer #1

Narrow sense heritability = h2 = Va/Vp, variance due to the effects of many alleles all added together (additive genetic variance = Va).

Vp = Va + Vd + Vi + Ve

Additive genetic variance = Va

Dominance genetic variance = Vd

Epistatic variance = Vi

Environmental variance = Ve

h2 = Va/Vp

Vp x h2 = Va

Vp = Va/h2

Va + Vd + Vi + Ve = Va/h2

Vd = (Va/h2) - Va - Vi - Ve

The higher the narrow sense heritability value (h2), the lower the dominance genetic variance, therefore trait A and B would be the the ones with the largest dominance genetic variance.

Answer: Traits A and B